OpenStudy (anonymous):

Please help. Will give medal and fan(: Add the following algebraic expression. 13/x-5 + 3/x+5

OpenStudy (whpalmer4):

\[\frac{13}{x-5}+\frac{3}{x+5}\]You need to establish a common denominator. The easy way to do this would be to multiply the left hand fraction by \(\dfrac{x+5}{x+5}\) and the right hand fraction by \(\dfrac{x-5}{x-5}\). Both fractions are equal to 1, so you aren't changing the value of anything, just putting it in a more convenient form.

OpenStudy (whpalmer4):

That gives you: \[\frac{13}{x-5}*\frac{x+5}{x+5} + \frac{3}{x+5}*\frac{x-5}{x-5} = \frac{13(x+5)+ 3(x-5)}{(x+5)(x-5)}\]

OpenStudy (whpalmer4):

Now distribute the stuff in the numerator and simplify it.

OpenStudy (anonymous):

16x + 50 would this be the top part?

OpenStudy (whpalmer4):

It would be indeed. Good job!

OpenStudy (anonymous):

(: your really good at teaching, i suck at fractions and look at that.

OpenStudy (whpalmer4):

Thanks, I try! It's my firm belief that much of what makes people think "math is hard" is just that they haven't encountered the right explanation yet. In a classroom with 30 students, the teacher can't really tailor an explanation for each students, but working one-to-one, I have the luxury of trying different explanations until we find one that works.

OpenStudy (whpalmer4):

Make that "each student" — I ken rilly speek the Eenglish properlee, honest! ;-)

OpenStudy (anonymous):

:D I completely agree in school the teachers aren't focusing on each kid. everybody learns different and i defiantly have trouble in math, but the way you were explaining made stuff click. haha

OpenStudy (anonymous):

Are you a teacher?

OpenStudy (whpalmer4):

Nah, I'm an engineer with a 10-year-old kid who likes to ask questions :-)

OpenStudy (anonymous):

im surprised you should really check out a teaching job, you would do wonders.

OpenStudy (whpalmer4):

Well, a few problems with that: 1) don't get to pick the students 2) the working conditions are usually awful, as is the pay My sister is a school teacher (not math, she's hopeless with math). She's always complaining about the job. I'm content to help out in my son's classroom on occasion, and here on OpenStudy.

OpenStudy (anonymous):

your a life saver on here. haha usually people just throw me the answer. thanks to you im starting to understand,your kid is lucky my parents are also hopeless at math.

OpenStudy (whpalmer4):

Well, if you get stuck, just tag me and I'll try to respond the next time I'm on. Even if someone else has answered, I'll check it over and add any comments I have (sometimes that consists of saying "uh, that's wrong!", unfortunately!)

OpenStudy (anonymous):

Deal.c: